# Supernovae in 2023 (review): breakthroughs by late observations

### 03 Jan 2024 v2

@@ 2 The point-symmetric morphology of SNRs
The study of SNRs (their properties, their progenitors, their interaction with the ISM, particle acceleration, and more) is a subject by itself that continued to be a major topic in 2023 (e.g., Araya 2023; Bakış et al. 2023; Ball et al. 2023; Bamba et al. 2023; Banovetz et al. 2023; Bozzetto et al. 2023; Braun et al. 2023; Bykov et al. 2023; Caldwell & Raymond 2023; Cao et al. 2023; Chambery et al. 2023; Deng et al. 2023; Dokara et al. 2023; Domček et al. 2023; Duvidovich & Petriella 2023; Eagle et al. 2023; Ellien, Greco, & Vink 2023; Ferrazzoli et al. 2023; Fujishige et al. 2023; Ghavamian et al. 2023; Godinaud et al. 2023; Himono et al. 2023; Holland-Ashford, Slane, & Long 2023; Katsuda 2023; Khabibullin et al. 2023; Kim, Koo, & Onaka 2023; Koplitz et al. 2023; Makarenko et al. 2023; Meyer et al. 2023; Narita et al. 2023; Nguyen-Dang et al. 2023; Ranasinghe & Leahy 2023; Rho et al. 2023a, b; Sankrit, Blair, & Raymond 2023; Sano et al. 2023; Seward & Points 2023; Shanahan et al. 2023; Sharma et al. 2023a; Sinitsyna & Sinitsyna 2023; Suherli et al. 2023; Sushch & Brose 2023; Suzuki et al. 2023; Vasiliev & Shchekinov 2023; Winkler, Long, & Blair 2023; Winkler et al. 2023; Xiao et al. 2023; Xin & Guo 2023; Yamauchi & Pannuti 2023; Yeung, Bamba, & Sano 2023; Zhang et al. 2023b, c; Zhang & Xin 2023; Zhong et al. 2023; Zhou et al. 2023; Alsaberi et al. 2024; López-Caraballo et al. 2024; Reyes-Iturbide et al. 2024). I will focus only on the morphologies of some SNRs.
@@ 3.1 Neutrino-driven versus jet-driven explosions
I start by reviewing some (but not all) studies of CCSNe that are relevant to the quest for the explosion mechanism. In comparing theory with observations, one must bear in mind the large diversity of CCSNe, as studies in 2023 continued to show (e.g., Agudo et al. 2023; Ailawadhi et al. 2023; Aryan et al. 2023; Ben-Ami et al. 2023; Blanchard et al. 2023; Bostroem et al. 2023a, c; Chen et al. 2023b; Corsi et al. 2023; Desai et al. 2023a; Dong et al. 2023b; Dorn-Wallenstein, Neugent, & Levesque 2023; Gangopadhyay et al. 2023; Holmbo et al. 2023; Irani et al. 2023; Jeena, Banerjee, & Heger 2023; Jiang & Liu 2023; Jin, Yoon, & Blinnikov 2023; Kilpatrick et al. 2023b; Könyves-Tóth & Seli 2023; Kuncarayakti et al. 2023; Li, Zhong, & Dai 2023; Lin, Zhang, & Zhang 2023a; Matsuoka et al. 2023; Moore et al. 2023; Moriya et al. 2023c; Park, Yoon, & Blinnikov 2023; Pessi et al. 2023b; Pessi, Prieto, & Dessart 2023; Pursiainen et al. 2023b; Rodríguez, Maoz, & Nakar 2023; Scully, Matzner, & Yalinewich 2023; Shivkumar et al. 2023; Sit et al. 2023; Srinivasaragavan et al. 2023; Stritzinger et al. 2023a, b; Subrayan et al. 2023; Sun, Maund, & Crowther 2023; Teja et al. 2023; Utrobin & Chugai 2023; Van Dyk et al. 2023a; Vasylyev et al. 2023a; Vazquez et al. 2023; Wang et al. 2023a, b; Wen et al. 2023; Williamson et al. 2023; Yan et al. 2023; Zhu et al. 2023; Bjornsson 2023).
I refer here to two explosion mechanisms, the delayed neutrino explosion mechanism and the JJEM. There are many new studies in 2023 of different aspects of the delayed neutrino explosion mechanism (e.g., Aguilera-Dena et al. 2023; Barker, O’Connor, & Couch 2023; Boccioli et al. 2023; Burrows, Vartanyan, & Wang 2023; Gogilashvili, Murphy, & Miller 2023; Navó et al. 2023; Orlando 2023; Powell et al. 2023; Sieverding, Kresse, & Janka 2023; Sumiyoshi et al. 2023; van Baal et al. 2023; Wang & Burrows 2023a, b, 2024; Zha et al. 2023; Maunder et al. 2024; Ghodla & Eldridge 2023). Varma, Müller, & Schneider (2023), for example, find with their 3D simulations that magnetic fields help drive earlier and more energetic explosions (see also Matsumoto, Takiwaki, & Kotake 2024). There is a lower number of new studies of the JJEM (e.g., Antoni & Quataert 2023; Shishkin & Soker 2023; Soker 2023e), including the operation of the JJEM in electron capture supernovae (Wang, Shishkin & Soker, 2024), that potentially can take place for stars with an initial mass of $8.5M_{\odot}\lesssim M_{\rm ZAMS}\lesssim 9.2M_{\odot}$ (e.g., Limongi et al. 2023). I will devote Section 3.2 to describe new results that I consider to strongly support the JJEM, and are toward a breakthrough in answering the question of the CCSN explosion mechanism. I therefore hope this review will motivate more studies of the JJEM.
@@
Some CCSNe require a central engine to operate after the explosion itself. Many studies attribute the extra energy to a rapidly rotating magnetized NS, i.e., a magnetar (e.g., Acharyya et al. 2023; Bai et al. 2023; Chen et al. 2023c; Gkini et al. 2023; Hu et al. 2023b; Omand & Sarin 2023; Prasanna et al. 2023; Tinyanont et al. 2023; Wang et al. 2023). As I pointed out in the past, the formation process of an energetic magnetar is accompanied by the launching of even more energetic jets, during and/or after the explosion. The majority of magnetar studies continue to ignore the role of jets. The problems of fitting the lightcurves of superluminous CCSNe with magnetars continue with 2023 papers. In Soker (2022) I analyzed the paper by Chen et al. (2023d) who fitted the lightcurves of 70 hydrogen-poor superluminous CCSNe with magnetars, about a third of these have an initial spin period of $P_{\rm mag}\leq 0.002{~{}\rm s}$. However, such rapidly rotating magnetars are likely to be spun up during the last phase of their formation by an accretion disk/belt that launches jets. I argued there that even the slower magnetars in their fitting most likely launched jets at the explosion. In seven out of the 14 CCSNe they fit with ejecta-CSM interaction, the explosion energy is much larger than what the delayed neutrino explosion mechanism can supply, implying a jet-driven explosion. The claim that supernova studies that include energetic magnetars cannot ignore jets continues to hold in 2023. The same arguments hold for the magnetars with $P_{\rm mag}\leq 0.002{~{}\rm s}$ that West et al. (2023) and Poidevin et al. (2023) find to fit the hydrogen-poor superluminous SN 2020qlb and SN 2021fpl, respectively, for the number of short-period magnetars that Dong et al. (2023c) argue to power bumpy lightcurves, for those magnetars with $P_{\rm mag}\leq 0.002{~{}\rm s}$ that Gomez et al. (2023) and Hinkle, Shappee, & Tucker (2023) fit to hydrogen-poor superluminous CCSNe, and to the magnetar that Fiore et al. (2023) fit to SN 2019neq.
@@
Studies in 2023 continue the exploration of gravitational waves from CCSNe and their detectability (e.g., Bruel et al. 2023; Bugli et al. 2023; Casallas Lagos et al. 2023; Dálya et al. 2023; Heston et al. 2023; Hsieh et al. 2023; Iess et al. 2023; Jakobus et al. 2023; Mezzacappa et al. 2023; Mitra et al. 2023; Pajkos et al. 2023; Pastor-Marcos et al. 2023; Shibagaki et al. 2023; Szczepańczyk et al. 2023; Vartanyan et al. 2023; Villegas et al. 2023; Wolfe et al. 2023; Yuan et al. 2023), including from two cocoons (lobes) that are inflated by fixed-axis energetic jets related to CCSNe (Gottlieb et al., 2023) and by several jet-inflated pairs of lobes (cocoons) in the JJEM (Soker, 2023c). The detection of neutrinos from CCSNe continued to be a topic of interest in 2023 (e.g., Abbasi et al. 2023; Akaho, Nagakura, & Foglizzo 2023; Bendahman et al. 2023; Drago et al. 2023; Fiorillo et al. 2023; Lin et al. 2023b; Mukhopadhyay & Sen 2023; Prasanna, Thompson, & Hirata 2023; Vartanyan & Burrows 2023; Powell et al. 2023). The detection of gravitational waves or neutrinos is not at a stage that allows to support of one explosion mechanism or the other (for the study of CCSN diagnostics see, e.g., Fryer et al. 2023). Therefore, I will not touch on these topics. I will also not touch the subject of axion instability supernova (e.g., Mori et al. 2023).
@@
The presence of CSM around CCSN progenitors and the interaction of the CCSN ejecta with this CSM were major subjects in the 2023 supernova literature (e.g., Akashi, Michaelis, & Soker 2023; Andrews et al. 2023; Ben-Ami et al. 2023; Bietenholz et al. 2023; Bilinski et al. 2023; Bruch et al. 2023; Chugai & Utrobin 2023; Das et al. 2023; Davis et al. 2023; Dessart & Jacobson-Galán 2023; Dickinson et al. 2023; Ercolino et al. 2023; Hiramatsu et al. 2023a; Ishii et al. 2023; Khatami & Kasen 2023; Koo et al. 2023a; Lin et al. 2023; Maeda et al. 2023a, c; Matsuoka & Sawada 2023; Mauerhan et al. 2023; Moran et al. 2023; Moriya 2023; Pearson et al. 2023a; Pellegrino et al. 2023; Pessi et al. 2023a; Petruk et al. 2023; Pursiainen et al. 2023a; Salmaso et al. 2023; Sarmah et al. 2023; Shrestha et al. 2023; Takei et al. 2023; Tsuna, Murase, & Moriya 2023; Tsuna & Takei 2023; Brennan et al. 2023; Murase 2024; Sfaradi et al. 2023), including the role of binary interaction in expelling pre-explosion CSM (e.g., Zenati et al. 2024). [^1] Soker (2013) suggested (based on earlier binary models of similar types of transients) that a sudden pre-CCSN swelling may trigger asymmetric mass loss in binary systems. Five months later Smith & Arnett (2014) repeated this suggestion. Later papers (e.g., Smith 2017; Smith et al. 2023; Bilinski et al. 2023; Mauerhan et al. 2023; Smith et al. 2023) wrongly and consistently have been attributing this suggestion to Smith & Arnett (2014). The ejecta-CSM interaction has become a major topic in relation to the type II SN 2023ixf (e.g., Bersten et al. 2023; Chandra et al. 2023; Guetta et al. 2023; Hiramatsu et al. 2023b; Hosseinzadeh et al. 2023a; Koenig 2023; Liu et al. 2023a; Martinez et al. 2023; Niu et al. 2023; Pledger & Shara 2023; Qin et al. 2023; Sarmah 2023; Sgro et al. 2023; Teja et al. 2023; Van Dyk et al. 2023b; Vasylyev et al. 2023b; Xiang et al. 2023a; Yamanaka, Fujii, & Nagayama 2023; Zimmerman et al. 2023).
@@
Studies found that the ejecta of SN 2023ixf interacted with a CSM that was extended up to a radius of $R_{\rm CSM}\simeq 20-50{~{}\rm AU}$ around SN 2023ixf (e.g., Berger et al. 2023; Bostroem et al. 2023b; Grefenstette et al. 2023; Jacobson-Galán et al. 2023; Kilpatrick et al. 2023a; Teja et al. 2023; Smith et al. 2023; Zhang et al. 2023a; Li et al. 2023). In principle, a close CSM might result from an enhanced mass loss rate shortly before the explosion, most likely accompanied by an outburst (eruption; e.g., Tsuna, Takei, & Shigeyama 2023) that might be triggered by activity in the core (e.g. Cohen & Soker 2024). However, several studies find no evidence for pre-explosion outbursts in the years before the explosion of SN 2023ixf (e.g., Dong et al. 2023a; Flinner et al. 2023; Neustadt, Kochanek, & Smith 2023; Jencson et al. 2023; Soraisam et al. 2023; Ransome et al. 2024). This challenges any model where the massive CSM is attributed to impulsive mass ejection years before the explosion of SN 2023ixf. The progenitor of SN 2023ixf did have large amplitude pulsations (e.g., Kilpatrick et al. 2023a; Soraisam et al. 2023). These pre-explosion properties led to the suggestion that instead of a pre-explosion impulsive ejection of a massive CSM, the progenitor of SN 2023ixf maintained an extended zone above its photosphere of rising and falling dense clumps, i.e., the effervescent zone model (Soker, 2023a).
@@
Obergaulinger & Reichert (2023) study nucleosynthesis in fixed-axis jets that power some energetic CCSNe, and find that they have enhanced synthesis of iron-group elements. Leung, Nomoto, & Suzuki (2023) and Leung & Nomoto (2024) study nucleosynthesis in CCSNe driven by fixed-axis jets, including collapsar where the remnant is a black hole. Wanajo (2023) studies nucleosynthesis in neutrino-heated ejecta and neutrino-driven winds, and Friedland, Mukhopadhyay, & Patwardhan (2024) and Psaltis et al. (2023) study the $\nu p$-process in neutrino-driven outflows in CCSNe. These papers do not refer to jets. In general, r-process nucleosynthesis in CCSNe was not a major topic in 2023. Anand et al. (2023), for example, study the lightcurves of broad-line stripped-envelope CCSNe and find no evidence for r-process nucleosynthesis. Nucleosynthesis can be used to study CCSNe. Hoppe et al. (2023) and Liu et al. (2024) present interesting studies of using solar-system isotopes ratio to constrain the stellar properties at the explosion.

### 15 Mar 2024 v3

@@ 2 The point-symmetric morphology of SNRs
The study of SNRs (their properties, their progenitors, their interaction with the ISM, particle acceleration, and more) is a subject by itself that continued to be a major topic in 2023 (e.g., Araya 2023; Bakış et al. 2023; Ball et al. 2023; Bamba et al. 2023; Banovetz et al. 2023; Bozzetto et al. 2023; Braun et al. 2023; Bykov et al. 2023; Caldwell & Raymond 2023; Cao et al. 2023; Chambery et al. 2023; Deng et al. 2023; Dokara et al. 2023; Domček et al. 2023; Duvidovich & Petriella 2023; Eagle et al. 2023; Ellien, Greco, & Vink 2023; Ferrazzoli et al. 2023; Fujishige et al. 2023; Ghavamian et al. 2023; Godinaud et al. 2023; Himono et al. 2023; Holland-Ashford, Slane, & Long 2023; Katsuda 2023; Khabibullin et al. 2023; Kim, Koo, & Onaka 2023; Koplitz et al. 2023; Makarenko et al. 2023; Meyer et al. 2023; Narita et al. 2023; Nguyen-Dang et al. 2023; Ranasinghe & Leahy 2023; Rho et al. 2023a, b; Sankrit, Blair, & Raymond 2023; Sano et al. 2023; Seward & Points 2023; Shanahan et al. 2023; Sharma et al. 2023a; Sinitsyna & Sinitsyna 2023; Suherli et al. 2023; Sushch & Brose 2023; Suzuki et al. 2023; Vasiliev & Shchekinov 2023; Winkler, Long, & Blair 2023; Winkler et al. 2023; Xiao et al. 2023; Xin & Guo 2023; Yamauchi & Pannuti 2023; Yeung, Bamba, & Sano 2023; Zhang et al. 2023b, c; Zhang & Xin 2023; Zhong et al. 2023; Zhou et al. 2023; Alsaberi et al. 2024a, b; Arias et al. 2024; Bao et al. 2024; Fesen et al. 2024; Khabibullin et al. 2024; Kirchschlager et al. 2024; López-Caraballo et al. 2024; Luo, Tang, & Mo 2024; Mantovanini et al. 2024; Michailidis et al. 2024a, b; Milisavljevic et al. 2024; Picquenot et al. 2024; Reyes-Iturbide et al. 2024; Reynolds & Borkowski 2024; Romano, Behrendt, & Burkert 2024; Sato et al. 2024; Vink et al. 2024; Zhang et al. 2024; Zhou et al. 2024). I will focus only on the morphologies of some SNRs.
@@ 3.1 Neutrino-driven versus jet-driven explosions
I start by reviewing some (but not all) studies of CCSNe that are relevant to the quest for the explosion mechanism. In comparing theory with observations, one must bear in mind the large diversity of CCSNe, as studies in 2023 continued to show (e.g., Agudo et al. 2023; Ailawadhi et al. 2023; Aryan et al. 2023; Ben-Ami et al. 2023; Blanchard et al. 2023; Bostroem et al. 2023a, c; Chen et al. 2023b; Corsi et al. 2023; Desai et al. 2023a; Dong et al. 2023b; Dorn-Wallenstein, Neugent, & Levesque 2023; Gangopadhyay et al. 2023; Holmbo et al. 2023; Irani et al. 2023; Jeena, Banerjee, & Heger 2023; Jiang & Liu 2023; Jin, Yoon, & Blinnikov 2023; Kilpatrick et al. 2023b; Könyves-Tóth & Seli 2023; Kuncarayakti et al. 2023; Li, Zhong, & Dai 2023; Lin, Zhang, & Zhang 2023a; Matsuoka et al. 2023; Moore et al. 2023; Moriya et al. 2023c; Park, Yoon, & Blinnikov 2023; Pessi et al. 2023b; Pessi, Prieto, & Dessart 2023; Pursiainen et al. 2023b; Rodríguez, Maoz, & Nakar 2023; Scully, Matzner, & Yalinewich 2023; Shivkumar et al. 2023; Sit et al. 2023; Srinivasaragavan et al. 2023; Stritzinger et al. 2023a, b; Subrayan et al. 2023; Sun, Maund, & Crowther 2023; Teja et al. 2023; Utrobin & Chugai 2023; Van Dyk et al. 2023a; Vasylyev et al. 2023a; Vazquez et al. 2023; Wang et al. 2023a, b; Wen et al. 2023; Williamson et al. 2023; Yan et al. 2023; Zhu et al. 2023; Bjornsson 2023; Brennan et al. 2024b; Burrows, Wang, & Vartanyan 2024; Chugai & Utrobin 2024; Das et al. 2024; Dastidar et al. 2024; Dessart et al. 2024; Dessart, Hillier, & Leonard 2024b; Elias-Rosa et al. 2024; Kangas et al. 2024; Lin et al. 2024; Sato et al. 2024; Shahbandeh et al. 2024; Sharma et al. 2024).
I refer here to two explosion mechanisms: the delayed neutrino explosion mechanism and the JJEM. There are many new studies in 2023 of different aspects of the delayed neutrino explosion mechanism (e.g., Aguilera-Dena et al. 2023; Barker, O’Connor, & Couch 2023; Boccioli et al. 2023; Burrows, Vartanyan, & Wang 2023; Gogilashvili, Murphy, & Miller 2023; Navó et al. 2023; Orlando 2023; Powell et al. 2023; Sieverding, Kresse, & Janka 2023; Sumiyoshi et al. 2023; van Baal et al. 2023; Wang & Burrows 2023a, b; Zha et al. 2023; Andresen et al. 2024; Ghodla & Eldridge 2023; Janka & Kresse 2024; Maunder et al. 2024; Schneider, Podsiadlowski, & Laplace 2024; Wang & Burrows 2024). Varma, Müller, & Schneider (2023), for example, find with their 3D simulations that magnetic fields help drive earlier and more energetic explosions (see also Matsumoto, Takiwaki, & Kotake 2024). There is a lower number of new studies of the JJEM (e.g., Antoni & Quataert 2023; Bear & Soker 2023; Shishkin & Soker 2023; Soker 2023e), including the operation of the JJEM in electron capture supernovae (Wang, Shishkin, & Soker, 2024), that potentially can take place for stars with an initial mass of $8.5M_{\odot}\lesssim M_{\rm ZAMS}\lesssim 9.2M_{\odot}$ (e.g., Guo et al. 2024; Limongi et al. 2023). I will devote Section 3.2 to describe new results that I consider to strongly support the JJEM, and are toward a breakthrough in answering the question of the CCSN explosion mechanism. I therefore hope this review will motivate more studies of the JJEM.
@@
Some CCSNe require a central engine to operate after the explosion itself (e.g., Matsumoto, Metzger, & Goldberg 2024). Many studies attribute the extra energy to a rapidly rotating magnetized NS, i.e., a magnetar (e.g., Acharyya et al. 2023; Bai et al. 2023; Chen et al. 2023c; Gkini et al. 2023; Hu et al. 2023b; Omand & Sarin 2023; Prasanna et al. 2023; Tinyanont et al. 2023; Wang et al. 2023). As I pointed out in the past, the formation process of an energetic magnetar is accompanied by the launching of even more energetic jets, during and/or after the explosion. The majority of magnetar studies continue to ignore the role of jets. The problems of fitting the lightcurves of superluminous CCSNe with magnetars continue with 2023 papers. In Soker (2022) I analyzed the paper by Chen et al. (2023d) who fitted the lightcurves of 70 hydrogen-poor superluminous CCSNe with magnetars, about a third of these have an initial spin period of $P_{\rm mag}\leq 0.002{~{}\rm s}$. However, such rapidly rotating magnetars are likely to be spun up during the last phase of their formation by an accretion disk/belt that launches jets. I argued there that even the slower magnetars in their fitting most likely launched jets at the explosion. In seven out of the 14 CCSNe they fit with ejecta-CSM interaction, the explosion energy is much larger than what the delayed neutrino explosion mechanism can supply, implying a jet-driven explosion. The claim that supernova studies that include energetic magnetars cannot ignore jets continues to hold in 2023. The same arguments hold for the magnetars with $P_{\rm mag}\leq 0.002{~{}\rm s}$ that West et al. (2023) and Poidevin et al. (2023) find to fit the hydrogen-poor superluminous SN 2020qlb and SN 2021fpl, respectively, for the number of short-period magnetars that Dong et al. (2023c) argue to power bumpy lightcurves, for those magnetars with $P_{\rm mag}\leq 0.002{~{}\rm s}$ that Gomez et al. (2023) and Hinkle, Shappee, & Tucker (2023) fit to hydrogen-poor superluminous CCSNe, and to the magnetar that Fiore et al. (2023) fit to SN 2019neq.
@@
Studies in 2023 continue the exploration of gravitational waves from CCSNe and their detectability (e.g., Bruel et al. 2023; Bugli et al. 2023; Casallas Lagos et al. 2023; Dálya et al. 2023; Heston et al. 2023; Hsieh et al. 2023; Iess et al. 2023; Jakobus et al. 2023; Mezzacappa et al. 2023; Mitra et al. 2023; Pajkos et al. 2023; Pastor-Marcos et al. 2023; Shibagaki et al. 2023; Szczepańczyk et al. 2023; Vartanyan et al. 2023; Villegas et al. 2023; Wolfe et al. 2023; Yuan et al. 2023; Nunes et al. 2024; Zha, Eggenberger Andersen, & O’Connor 2024a), including from two cocoons (lobes) that are inflated by fixed-axis energetic jets related to CCSNe (Gottlieb et al., 2023) and by several jet-inflated pairs of lobes (cocoons) in the JJEM (Soker, 2023c). The detection of neutrinos from CCSNe continued to be a topic of interest in 2023 (e.g., Abbasi et al. 2023; Akaho, Nagakura, & Foglizzo 2023; Bendahman et al. 2023; Drago et al. 2023; Fiorillo et al. 2023; Lin et al. 2023b; Mukhopadhyay & Sen 2023; Pagliaroli & Ternes 2024; Prasanna, Thompson, & Hirata 2023; Vartanyan & Burrows 2023; Ashida 2024; Powell et al. 2023; Saez et al. 2024). The detection of gravitational waves or neutrinos is not at a stage that allows the support of one explosion mechanism or the other (for the study of CCSN diagnostics, see, e.g., Fryer et al. 2023). Therefore, I will not touch on these topics. I will also not touch the subject of axion instability supernova (e.g., Mori et al. 2023).
@@
The presence of CSM around CCSN progenitors and the interaction of the CCSN ejecta with this CSM were major subjects in the 2023 supernova literature (e.g., Akashi, Michaelis, & Soker 2023; Andrews et al. 2023; Ben-Ami et al. 2023; Bietenholz et al. 2023; Bilinski et al. 2023; Bruch et al. 2023; Chugai & Utrobin 2023; Das et al. 2023; Davis et al. 2023; Dessart & Jacobson-Galán 2023; Dickinson et al. 2023; Ercolino et al. 2023; Hiramatsu et al. 2023a; Ishii et al. 2023; Khatami & Kasen 2023; Koo et al. 2023a; Lin et al. 2023; Maeda et al. 2023a, c; Matsuoka & Sawada 2023; Mauerhan et al. 2023; Moran et al. 2023; Moriya 2023; Pearson et al. 2023a; Pellegrino et al. 2023; Pessi et al. 2023a; Petruk et al. 2023; Pursiainen et al. 2023a; Salmaso et al. 2023; Sarmah et al. 2023; Shrestha et al. 2023; Takei et al. 2023; Tsuna, Murase, & Moriya 2023; Tsuna & Takei 2023; Brennan et al. 2024a; Ferrari et al. 2024; Jacobson-Galán et al. 2024; Murase 2024; Meza Retamal et al. 2024; Murai et al. 2024; Sfaradi et al. 2024), including the role of binary interaction in expelling pre-explosion CSM (e.g., Zenati et al. 2024). [^1] Soker (2013) suggested (based on earlier binary models of similar types of transients) that a sudden pre-CCSN swelling may trigger asymmetric mass loss in binary systems. Five months later Smith & Arnett (2014) repeated this suggestion. Later papers (e.g., Smith 2017; Smith et al. 2023; Bilinski et al. 2023; Mauerhan et al. 2023; Smith et al. 2023) wrongly and consistently have been attributing this suggestion to Smith & Arnett (2014). The ejecta-CSM interaction has become a major topic in relation to the type II SN 2023ixf (e.g., Bersten et al. 2023; Chandra et al. 2023; Guetta et al. 2023; Hiramatsu et al. 2023b; Hosseinzadeh et al. 2023a; Koenig 2023; Liu et al. 2023a; Martinez et al. 2023; Niu et al. 2023; Pledger & Shara 2023; Qin et al. 2023; Sarmah 2023; Sgro et al. 2023; Teja et al. 2023; Van Dyk et al. 2023b; Vasylyev et al. 2023b; Xiang et al. 2023a; Yamanaka, Fujii, & Nagayama 2023; Zimmerman et al. 2023).
@@
Studies found that the ejecta of SN 2023ixf interacted with a CSM that was extended up to a radius of $R_{\rm CSM}\simeq 20-50{~{}\rm AU}$ around SN 2023ixf (e.g., Berger et al. 2023; Bostroem et al. 2023b; Grefenstette et al. 2023; Jacobson-Galán et al. 2023; Kilpatrick et al. 2023a; Teja et al. 2023; Smith et al. 2023; Zhang et al. 2023a; Li et al. 2023). In principle, a close CSM might result from an enhanced mass loss rate shortly before the explosion, most likely accompanied by an outburst (eruption; e.g., Tsuna, Takei, & Shigeyama 2023; Tsuna et al. 2024) that might be triggered by activity in the core (e.g. Cohen & Soker 2024). However, several studies find no evidence for pre-explosion outbursts in the years before the explosion of SN 2023ixf (e.g., Dong et al. 2023a; Flinner et al. 2023; Neustadt, Kochanek, & Smith 2023; Jencson et al. 2023; Soraisam et al. 2023; Ransome et al. 2024). This challenges any model where the massive CSM is attributed to impulsive mass ejection years before the explosion of SN 2023ixf. The progenitor of SN 2023ixf did have large amplitude pulsations (e.g., Kilpatrick et al. 2023a; Soraisam et al. 2023). These pre-explosion properties led to the suggestion that instead of a pre-explosion impulsive ejection of a massive CSM, the progenitor of SN 2023ixf maintained an extended zone above its photosphere of rising and falling dense clumps, i.e., the effervescent zone model (Soker, 2023a).
@@
Obergaulinger & Reichert (2023) study nucleosynthesis in fixed-axis jets that power some energetic CCSNe, and find that they have enhanced synthesis of iron-group elements. Reichert et al. (2024) find that when magnetic fields are included in the fixed-axis simulations, nucleosynthesis can yield up to the third peak of the r-process (but see Zha, Müller, & Powell 2024b). Leung, Nomoto, & Suzuki (2023) and Leung & Nomoto (2024) study nucleosynthesis in CCSNe driven by fixed-axis jets, including collapsar where the remnant is a black hole. Wanajo (2023) studies nucleosynthesis in neutrino-heated ejecta and neutrino-driven winds (see also Prasanna, Coleman, & Thompson (2024)), and Friedland, Mukhopadhyay, & Patwardhan (2024) and Psaltis et al. (2023) study the $\nu p$-process in neutrino-driven outflows in CCSNe. These papers do not refer to jets. In general, r-process nucleosynthesis in CCSNe was not a major topic in 2023. Anand et al. (2023), for example, study the lightcurves of broad-line stripped-envelope CCSNe and find no evidence for r-process nucleosynthesis. Patel et al. (2024) study the effect of r-process nucleosynthesis on the lightcurve of hydrogen-rich CCSNe. Nucleosynthesis can be used to study CCSNe. Hoppe et al. (2023) and Liu et al. (2024) present interesting studies of using solar-system isotopes ratio to constrain the stellar properties at the explosion.